FEL Upgrade Project Weekly Brief September 9-13, 2002



Highlights:
This week all the quadrupole magnet/girder assemblies for the recirculation arc re-injection sector were installed in the FEL Upgrade.

Work continues at a fast pace to complete the installation of the injector/gun hardware in preparation for start-up of commissioning activities next month.

The proceedings of the January 2002 FEL Users/LPC Workshop are posted at:
http://www.jlab.org/FEL/LPC/mtgs_wkshps.html.
Please mark your calendars for the next LPC Workshop to be held January 23-24, 2003.

Management:
Several members of the scientific and engineering staff attended the International FEL Conference held at Argonne National Labs this week.  Papers that were presented by the teams and our collaborators at AES, Inc and the Naval post Graduate School will be added to the FEL bibliography contained in the FEL website: http://www.jlab.org/FEL and the JLab Publication website: http://www.jlab.org/div_dept/admin/publications/02pub.html.

We thank the Office of Naval Research for their transfer of the last component of FY02 funds ($350k) for support of FEL Upgrade activities.

The paperwork was prepared for the required DOC export control declarations on the 1 kW Demo FEL wiggler, which be loaned to the Daresbury Laboratory (UK) for their planned 4GLS advanced light source project. Copies of this paperwork will be forwarded to the ONR Program Office.  The previously prepared loan agreement has already been sent to ONR.

WBS 4 (Injector):
The instrument flange drawing was signed off and submitted to the shop.  The rear bellows/drive and the rear corona shield drawings are in sign off which are the last gun drawings needed to complete the gun.  The gun oven was sent out for required modifications.  Modifications were made to the gun stand to correct an interference problem.  The ground end internal corona ring is being polished and nearly complete.  Assembly of the light box has been started.

Gun HVPS - The large toroid was aligned, the resistor mounting bracket made and welded onto the toroid by the Machine Shop, and the toroid reinstalled back in the tank this week.  The HVPS is ready for final wiring and testing.  A system diagram is about half done for the wiring and the interface to the PSS & MPS.

WBS 6 (RF):
Quarter HVPS - Wiring and upgrading of the HVPS continued this week.  Interlock tests are planned for early next week.  All of the RF couplers have been calibrated and reinstalled.  The LCW manifolds for the Circulator, RF Load, and HOM filter have been plumbed and leak tested.

Zone 1 - The waveguides were reconnected to the Quarter this week.

Zone 4 - The half height waveguides were completed by the Machine Shop and brought over to the FEL.  They will be installed when people are available.

WBS 8 (Instrumentation):
We are continuing investigating interest for commercial engineering firms to bid the regulator/DSP board design and assembly for a new trim power supply.  There are a number of firms that have excess capacity since the down-turn of the telecom market and are hungry.  We are pursuing purchasing a 60 amp current transducer from Danfisik as an OEM part.  They are producing ~200/month for MRI shim magnet power supply regulators.  This device would have 50 - 100X better performance (stability/reproducability) than the LEM module used in the Fermi switcher at a lower cost.

The schematic and plots for the Acromag fanout module have been completed, reviewed and awaiting sign off. Progress has continued for the design of the HV interlock chassis and fabrication has begun.  The HV interlock board is near completion.  Cable tray was hung above the SF6 tank and cabling for the HV has begun.

Plots for the GPIM fanout board have reviewed and ready for sign off. Fabrication of the Stepper Motor breakout chassis has begun.  Two (2) Beamviewer Canbus Interface modules have completed and turned over to A. Grippo for testing.

BLMs - The F0150 rev A (rear cards) are due back this coming Monday (9-16).  All parts are in house and ready for assembly.  The F0150 pcbs (front cards) are proceeding well thru the testing phase.  We should have our first block of 12 channels of BLM ready by October 1st.
DLPC - The CO305 carrier board pcb layout and schematic are finished and the package has been sent out for production on a 5 day turn.  It's implementation is anticipated by Oct. 1st as well.
MPS - PCB layout and schematics were checked and revised for the 6U breakout adapter (for the general purpose isolation module)
MISC - The photocathode scanner chassis is being recommissioned in Lab 5. Assistance in documenting the high voltage power supply design was provided as well.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
MEI is checking all the core gaps of the remaining six cores with our measuring fixture.
Injector Dipoles (DU/DV)
   The new hysteresis procedure spoiled the gradient and transverse gradient integral goodness of past
    measurements.  The integral is about 2.5% small, beyond the ability of the field clamps to correct.
    However, since the magnet is a wedge and there is lots of room between the coil and beam
    chamber, a movement of 1.7 cm of the fat side to the center gives the right integral.  Our
    measurements of the linearity of the transverse integral gradient show we are at ± 3 parts in 10,000.
    Dave's specification is ±1 part. We have fixes for that too.
Arc 180 Degree Dipoles (GY)
   GY Coil Potting fixture continues its preparation for potting.
   We were notified of the shipment of the extra copper for the fifth coil from Finland.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
   The GX cores were shipped on Wednesday to Wang NMR
   At Wang NMR, the first GQ coil was vacuum impregnated with epoxy.  No word yet on the cure
    and how it came out.
Quadrupoles
Trim Quad (QT)
   The first trim quad was fully measured and Tom Hiatt is getting the results to David Douglas in
    before going on with tests of the remaining magnets.
Sextupole (SF)
   DULY Research didn't finish the drawings as they had indicated they would.  We are pulling them
    as is and will deal with them in some other way.
   We are working on the design of the quick fix sextupoles.
Octupole (OT)
   The Octupole design continued at DULY Research.
Beam Line and Vacuum
   X- Chamber - Welding of transitions to chamber is in progress.
   Chicane Chamber - Wire EDM work for the transition piece has been awarded.  Welding of the
    first chicane chamber was started.  Machining of other chamber plates, tubes, and CFF's are on
    hold to allow higher priority parts to be fabricated.
  Y-Chamber - Plate machining is in progress.
   ARC Chambers - Master Machine has completed the fit up of the first chamber.  Full penetration
    fusion welding will begin tomorrow.  The chamber still requires finish machining of the weldment,
    cleaning, final fitting of the flanges, and leak checking.  For the other ARC chambers, the machining
    of the bevels is delayed due to a controller problem on the CNC machine.  The machine is currently
    down.  No estimate of the repair is available at this time.
   Design of girders for the regions around the wiggler were signed off and are in procurement.
   Girder and vacuum assembly continued
   Low Conductivity Water piping installation continued in the FEL enclosure
   The alignment crew continued stand placement.

WBS 11 (Optics):
The majority of the optics group attended the FEL Conference & User Workshop at Argonne.  Nevertheless, a significant amount of work occurred.  Two leadscrews were stripped of their original silver coating and sent to be sputter-coated with tungsten disulfide.  Speaking with colleagues at the conference, several said they have used tungsten disulfide in UHV, near optics, with no problems.

Water cooling was supplied to the 3" optic in the test and assembly stand, and the amount of noise quantified.  The mass of the yaw plate effectively dampened any noise on that axis.  For the pitch plane, the noise in the ~ 70 Hz region increased by ~ 15 dB.  When driven, the resonance frequencies shifted some as well, as was expected.  Given the small amount of additional noise, the feedback circuit is being modified to control it.

The design of the turning cassettes has made great progress, and we have ordered a large number of parts needed.  This was also true for the insertable mirror.

Other work:
We have chosen a vendor for the switch laser that will be used with the external cavity (aka the pulse stacker) for the FEL.